When Euclid dresses up for goth night, he has to choose a cloak, a shade of dark lipstick, and a pair of boots. He has 2 cloaks, 6 shades of dark lipstick, and 3 pair of boots. How many different choices of outfit can he create?
Answer by jim_thompson5910(35256)

There are 2*6*3 = 12*3 = 36 different outfits to choose from.